40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 69 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  14173   Tue Aug 21 09:16:23 2018 SteveUpdateWiki AP table layout 20180821

 

 

Attachment 1: 20180821.JPG
20180821.JPG
  14172   Tue Aug 21 03:09:59 2018 johannesOmnistructureDAQPanels for Acromag DAQ chassis

I expanded the previous panels to 6U height for the new DAQ chassis we're buying for the upgrade. I figure it's best if we stick to the modular design, so I'm showing a panel for 8 BNC connectors as an example. The front panel has 12 slots, the back has 10 plus power connectors, switches, and the ethernet plug.

I moved the power switch to the rear because it's a waste of space to put it in the front, and it's not like we're power cycling this thing all the time. Note that the unit only requires +24V (for general operation, +20V also does the trick, as is the situation for ETMX) and +15V (excitation field for the binary I/O modules). While these could fit into a single CONEC power connector, it's probably for the better if we don't make a version that supplies a large positive voltage where negative is expected, so I put in two CONEC plugs for +/- 15 and +/- 24.

I want to order 5-6 of these as soon as possible, so if anyone wants anything changed or sees a problem, please do tell!

Attachment 1: auxdaq_40m_6U_front.pdf
auxdaq_40m_6U_front.pdf
Attachment 2: auxdaq_40m_6U_rear.pdf
auxdaq_40m_6U_rear.pdf
Attachment 3: auxdaq_40m_6U_BNC.pdf
auxdaq_40m_6U_BNC.pdf
  14171   Mon Aug 20 15:16:39 2018 JonUpdateCDSRebooted c1lsc, slow machines

When I came in this morning no light was reaching the MC. One fast machine was dead, c1lsc, and a number of the slow machines: c1susaux, c1iool0, c1auxex, c1auxey, c1iscaux. Gautam walked me through reseting the slow machines manually and the fast machines via the reboot script. The computers are all back online and the MC is again able to lock.

  14170   Mon Aug 20 14:04:53 2018 johannesBureaucracyEquipment loanTwo C30642G PDs removed

EDIT: After discussing with Koji and checking the existing M2ISS PDs I put the two C30642G back and took two C30665GH (active diameter: 3mm) diodes. Only one of this type remains in storage.

I removed two C30642G photodiodes from the stash for the new M2ISS hardware and updated the wiki page accordingly.

  14169   Thu Aug 16 23:06:50 2018 gautamUpdateSUSAnother low noise bias path idea

I had a very fruitful discussion with Rich about this circuit today. He agreed with the overall architecture, but made the following suggestions (Attachment #1 shows the circuit with these suggestions incorporated):

  1. Use an Op27 instead of LT1128, as it is a more friendly part especially in these composite amplifier topologies. I confirmed that this doesn't affect the output voltage noise at 100 Hz, we will still limited by Johnson noise of the 15kohm series resistor.
  2. Take care of voltage distribution in the HV feedback path
    • I overlooked the fact that the passive filtering stage means that the DC current we can drive in the configuration I posted earlier is 150V / 25kohm = 6mA, whereas we'd like to be able to drive at least 10 mA, and probably want the ability to do 12 mA to leave some headroom.
    • At the same time, the feedback resistance shouldn't be too small such that the PA91 has to drive a significant current in the feedback path (we'd like to save that for the coil).
    • Changing the supply voltage of the PA91 from 150 V to 320 V, and changing the gain to x30 instead of x15 (by changing the feedback resistor from 14kohm to 29kohm), we can still drive 12 mA through the 25 kohms of series resistance. This will require getting new HV power supplies, as the KEPCO ones we have cannot handle these numbers.
    • The current limiting resistor is chosen to be 25ohms such that the PA91 is limited to ~26 mA. Of this, 300V / 30kohm ~ 10 mA will flow in the feedback path, which means under normal operation, 12 mA can safely flow through the coils.
    • Rich recommended using metal film resistors in the high voltage feedback path. However, these have a power rating, and also a voltage rating. By using 6x 5kohm resistors, the max power dissipated in each resistor is 50^2 / 5000 ~ 0.5 W, so we can get 0.6 W (or 1W?)  rated resistors which should do the job. I think the S102K or S104K series will do the job.
  3. Add a voltage monitoring capability.
    • This is implemented via a resistive voltage divider at the output of the PA91.
    • We can use an amplifier stage with whitening if necessary, but I think simply reading off the voltage across the terminating resistor in the ladder will be sufficient since this circuit will only have DC authority.
  4. Make a Spice model instead of LISO, to simulate transient effects.
    • I've made the model, investigating transients now.
  5. High voltage precautions:
    • When doing PCB layout, ensure the HV points have more than the default clearance. Rich recommends 100 mils.
    • Use a dual-diode (Schottky) as input protection for the Op27 (not yet implemented in Spice model).
    • Use a TVS diode for the moniotring circuit (not yet implemented in Spice model).
    • Make sure resistors and capacitors that see high voltage are rated with some safety margin.
  6. Consider using the PA95 (which Rich has tested and approves of) instead of the PA91. Does anyone have any opinions on this?

If all this sounds okay, I'd like to start making the PCB layout (with 5 such channels) so we can get a couple of trial boards and try this out in a couple of weeks. Per the current threat matrix and noises calculated, coil driver noise is still projected to be the main technical noise contribution in the 40m PonderSqueeze NB (more on this in a separate elog).

Quote:

Looks like this will do the job. I'm going to run this by Rich and get his input on whether this will work (this design has a few differences from Rich's design), and also on how to best protect from HV incidents.

Attachment 1: HVamp_schem.PDF
HVamp_schem.PDF
Attachment 2: Hvamp.zip
  14168   Thu Aug 16 14:48:14 2018 SteveUpdateVACwhy do we need a root pump?

Basic Pump Throughput Concepts

What is Pump Throughput?

The manufacturer of a vacuum pump supplies a chart for each pump showing pumping speed (volume in unit time) vs pressure. The example, for a fictitious pump, shows the pumping speed is substantially constant over a large pressure range.

By multiplying pumping speed by pressure at which that pumping speed occurs, we get a measure called pump throughput. We can tabulate those results, as shown in the table below, or plot them as a graph of pressure vs pump throughput. As is clear from the chart,  pump throughput (which might also be called mass flow) decreases proportionally with PRESSURE, at least over the pressure range where pumping speed is constant.

 

Pumping Speed Pressure Pressure x Pumping Speed
100 L/sec 10 torr 1000 torr.liter/sec
100 L/sec 1 torr 100 torr.liter/sec
100 L/sec 0.1 torr 10 torr.liter/sec
100 L/sec 0.01 torr

1 torr.liter/sec

The roughing pump speed actually will reach 0 l/s  at it's ultimate pressure performance.

Our roughing pump  pumping speed will slowly drop  as chamber pressure drops. Below 10 Torr this decrease is accelerated and bottoms out. This where the Root pump can help. See NASA evaluation of dry rough pumps...What is a root pump 

We have been operating succsessfully with a narrow margin. The danger is that the Maglev forline peaks at 4 Torr. This puts load on the small turbo TP2, TP3 &  large TP1

The temperature of these TP2 & 3  70 l/s drag turbos go up to 38 C and their  rotation speed slow to 45K rpm from 50K rpm because of the large volume 33,000 liters

Either high temp or low rotation speed of drag turbo or long time of overloading  can shut down the small turbo pumps......meaning: stop pumping, wait till they cool down

The manual gate valve installed helped to lower peak temp to 32C It just took too long.

We have been running with 2 external fans [one on TP1 & one on TP3]  for cooling and one aux drypump to help lowering the foreline pressure of TP2 & 3

The vacuum control upgrade should include adding root pump into the zero pumping speed range.

 

Atm1,   Pump speed chart:   TP1  turbo -red, root pump -blue and mechanical pump green. Note green color here representing an oily rotory pump. Our small drypumps [SH-100] typically run above 100 mTorr

                                           They are the forepump of TP2 & 3     Our pumpdown procedure: Oily Leybold rotory pumps ( with safety orifice 350 mT to atm ) rough to 500 mTorr

                                                                                                 Here we switch over to TP2 & 3 running at 50k RPM with drypumps SH-100 plus Aux Triscroll

                                                                                                 TP1- Maglev rotating full speed when V1 is opened at full volume at 500 mTorr 

                         History: the original design of the early 1990s had no dry scroll pumps. Oil free dry scrools replaced the oily forepumps of TP2 & TP3 in ~2002  at the cost of degrading the forline pressure somewhat.

                                     We had 2 temperature related Maglev failers in 2005 Aug 8 and 2006 April 5  Osaka advised us to use AUX fan to cool TP1  This helped. 

Atm2,   Wanted Root pump - Leybold EcoDry 65 plus  

Atm3,   Typical 8 hrs pumpdown from 2007 with TP2 & 3 

Atm4,   Last pumpdown zoomed in from 400 mT to 1mT with throttled gate  valve took 9 hrs  The foreline pressure of TP1 peaked at 290 mT, TP3 temperature peaked at 32C

            This technic is workable, but 9 hrs is too long.

Atm5,   The lowest pressure achived in  the 40m Vacuum Envelope 5e-7 Torr with pumps Maglev  ~300 l/s,  Cryo 1500 l/s  and 3 ion pumps of 500 l/s      [ in April 2002 at pumpdown 53 day 7 ] with annuloses at ~ 10 mTorr

Atm6,  Osaka TG390MCAB Throughput with screen ~300 L/s at 12 cfm backing pump

Attachment 1: PUMPSPEED_CHAR.pdf
PUMPSPEED_CHAR.pdf
Attachment 2: Leybold_Broschuere_8Seiten_EN_ANSICHT.pdf
Leybold_Broschuere_8Seiten_EN_ANSICHT.pdf
Attachment 3: pd65.jpg.png
pd65.jpg.png
Attachment 4: pd81completed.png
pd81completed.png
Attachment 5: best_.pdf
best_.pdf
Attachment 6: Osaka390.pdf
Osaka390.pdf
  14167   Thu Aug 16 07:50:28 2018 SteveUpdateVACpumpdown 81 at day 30

 

 

Attachment 1: pd81d30.png
pd81d30.png
  14166   Wed Aug 15 21:27:47 2018 gautamUpdateCDSCDS status update

Starting c1cal now, let's see if the other c1lsc FE models are affected at all... Moreover, since MC1 seems to be well-behaved, I'm going to restore the nominal eurocrate configuration (sans extender board) tomorrow.

  14165   Wed Aug 15 19:18:07 2018 gautamUpdateSUSAnother low noise bias path idea

I took another pass at this. Here is what I have now:

Attachment #1: Composite amplifier design to suppress voltage noise of PA91 at low frequencies.

Attachment #2: Transfer function from input to output.

Attachment #3: Top 5 voltage noise contributions for this topology.

Attachment #4: Current noises for this topology, comparison to current noise from fast path and slow DAC noise.

Attachment #5: LISO file for this topology.

Looks like this will do the job. I'm going to run this by Rich and get his input on whether this will work (this design has a few differences from Rich's design), and also on how to best protect from HV incidents.

Attachment 1: HV_Bias.pdf
HV_Bias.pdf
Attachment 2: HVamp_TF.pdf
HVamp_TF.pdf
Attachment 3: HVamp_noises.pdf
HVamp_noises.pdf
Attachment 4: currentNoises.pdf
currentNoises.pdf
Attachment 5: HVamp.fil.zip
  14164   Wed Aug 15 12:15:24 2018 gautamUpdateCOCMacroscopic SRC length for SR tuning

Summary:

It looks like we can have a stable SRC of length 4.044 m without getting any new mirrors, so this is an option to consider in the short-term.

Details:

  • The detailed calculations are in the git repo
  • The optical configuration is:
    • A single folding mirror approximately at the current SR3 location.
    • An SRM that is ~1.5m away from the above folding mirror. Which table the SRM goes on is still an open question, per the previous elog in this thread. 
  • The SRC length is chosen to be 4.044 m, which is what the modeling tells us we need for operating in the SR tuning instead of RSE.
  • Using this macroscopic length, I found that we could use a single folding mirror in the SRC, and that the existing (convex) G&H folding mirrors, which have a curvature of -700m, happily combine with our existing SRM (concave with a curvature of 142m) to give reasonable TMS and mode-matching to the arm cavity.
  • The existing SRM transmission of 10% may not be optimal but Kevin's calculations say we should still be able to see some squeezing (~0.8 dB) with this SRM.
  • Attachment #1 - corner plot of the distribution of TMS for the vertical and horizontal modes, as well as the mode-matching (averaged between the two modes) between the SRC and arm cavity.
  • Attachment #2 - histograms of the distributions of RoCs and lengths used to generate Attachment #1. The distributions were drawn from i.i.d Gaussian pdfs.

gautam 245pm: Koji pointed out that the G&H mirrors are coated for normal incidence, but looking at the measurement, it looks like the optic has T~75ppm at 45 degree incidence, which is maybe still okay. Alternatively, we could use the -600m SR3 as the single folding mirror in the SRC, at the expense of slightly reduced mode-matching between the arm cavity and SRC.

Attachment 1: SRC_MCMC_shortTerm.pdf
SRC_MCMC_shortTerm.pdf
Attachment 2: SRC_dists_shortTerm.pdf
SRC_dists_shortTerm.pdf
  14163   Tue Aug 14 23:14:24 2018 aaronUpdateOMCOMC scanning/aligning script

I made a script to scan the OMC length at each setpoint for the two TTs steering into the OMC. It is currently located on nodus at /users/aaron/OMC/scripts/OMC_lockScan.py.

I haven't tested it and used some ez.write syntax that I hadn't used before, so I'll have to double check it.

My other qualm is that I start with all PZTs set at 0, and step around alternative +/- values on each PZT at the same magnitude (for example, at some value of PZT1_PIT, PZT1_YAW, PZT2_PIT, I'll scan PZT2_YAW=1, then PZT2_YAW=-1, then PZT2_YAW=2). If there's strong hysteresis in the PZTs, this might be a problem.

  14162   Tue Aug 14 02:01:12 2018 gautamUpdateLSCDRMI locking - partial success

After tweaking the AS55 demod phase, SRM alignment, triggering settings, I got a few brief DRMI locks in tonight, I'm calling it a success (though this isn't really robust yet). The main things to do now are:

  • turn on all the boosts on the LSC loops - today I only managed to trigger the PRCL boost filters successfully without blowing up the lock.
  • measure all 3 loops, tweak gain as necessary.
  • Run some sensing lines, tune the demod phase.
  • The SRCL triggering is strange to me - SRCL loop is currently triggered on POP22_I, but the 2f1 buildup in the symmetric side does not say anything about the linearity of the SRCL error signal? Or are we just hoping the SRM is in the correct place and engaging the servo? Anyway, this setting seems to work but perhaps once the locking is more robust the triggering can be fixed.
  • do a quick NB - I expect the main change to be that the AS55_Q dark noise contribution would have gone up on account on the reduced amount of light at this port.

I think the main IFO characterization remaining to be done to determine the status of the IFO post vent is to measure the losses of the arm cavities. IMO, we will need to certainly fix the clipping at ETMY before we attempt some serious locking.

Attachment 1: DRMI.png
DRMI.png
  14161   Tue Aug 14 00:50:32 2018 gautamUpdateASSX arm ASS still not quite right?

While working on the single arm alignment, I noticed that today, i was able to get the X arm transmission back to ~1.22, and the GTRX to 0.52. These are closer to the values I remember from prior to the vent. Running the dither alignment promptly degrades both the green and IR transmissions. Since the pianosa SL7 upgrade, I can't use the sensoray to capture images, but to me, the spot looks a little off-center in Yaw on ETMX in this configuration, I've tried to show this in the phone grab (Atm #2). Maybe indicative of clipping somewhere upstream of ITMX?

Anyways, I'm pushing onwards for now, something to check out in the daytime.

Quote:

[koji, gautam]

After I effected the series resistance change for ETMX, the X arm ASS didn't work (i.e. IR transmission would degrade if the servo was run). Today, we succeeded in recovering a functional ASS servo yes.

We then tried to maximize GTRX using the PZT mirrors, but were only successful in reaching a maximum of 0.41. The value I remember from before the vent was 0.5, and indeed, with the IR alignment not quite optimized before we began this work, I saw GTRX of 0.48. But the IR dither servo signals indicate that the cavity axis may have shifted (spot position on the ITM, which is uncontrolled, seems to have drifred significantly, the Pitch signal doesn't stay on the StripTool scale anymore). So we may have to double check that the transmitted beam isn't falling off the GTRX DC PD.

Attachment 1: POXPOY.png
POXPOY.png
Attachment 2: IMG_7108.JPG
IMG_7108.JPG
  14160   Tue Aug 14 00:27:55 2018 gautamUpdateLSCLocking prep

In preparation for attempting some DRMI locking, I did the following:

  • Slow machine reboots for unresponsive c1psl, c1susaux and c1iscaux. The latter requried a manual burtrestore to recover the usual LSC PD whitening settings.
  • Shuttered AUX laser (which was on Standby anyways) - we should really install a remotely controllable shutter for this on the AS table.
  • Re-aligned PMC (half turn of knob in yaw, full turn in pitch) - IMC transmission 15,000cts ---> 15,600cts.
  • Squished sat. box cables at ITMX and ETMX.

Not related to this work, but I turned the Agilent NA off since we aren't using it immediately.

  14159   Mon Aug 13 20:21:10 2018 aaronUpdateOMCNew DAC for the OMC

[aaron, gautam]

We finished up making the new c1omc model  (screenshot attached).

The new channels are only four DAC for ASC into the OMC, and one DAC for the OMC length:

C1:OMC-ASC_PZT1_PIT
C1:OMC-ASC_PZT1_YAW
C1:OMC-ASC_PZT2_PIT
C1:OMC-ASC_PZT2_YAW
C1:OMC-PZT
 
The model compiles and we can change the channel values, so we are all set to do this OMC scan on the software side.
Attachment 1: c1omcSCREENSHOT.png
c1omcSCREENSHOT.png
  14158   Mon Aug 13 17:20:07 2018 aaronConfigurationUpgradeParts list for BHD

I've attached the diagram of what I mean.

There are a couple caveats and changes that would have to be made that are not included in this diagram, because they would be made on different tables.

  1. I moved MMT2, which means the other MMT optics would have to be adjusted to accomodate this. I checked out the configuration on the BS table and this seems doable with a small rotation of MMT1 and maybe PJ2.
  2. I don't know the best way to get the OMC REFL beam out... thoughts?
  3. This diagram is kind of crappy after my edits, someone should tell me how to avoid collapsing all layers when I open these layout diagrams in inkscape, because I ended up editing the layout in Acrobat instead, where the lack of object grouping caused a bunch of the optics to lose one or two lines along the way.
  4. I didn't include all beam paths explicitly but can if this looks like a good configuration.
  5. The optic that picks off the transmission through MMT2 will need to move a bit, but there was a clamp in the way; this should be a minor change.
  6. The optic just before the OMC needs to be moved up a bit.
  7. The optic after the signal OMMT should be changed to a PBS and translated a bit; this is where the LO and signal beams will combine

Gautam also had some questions about the BHD/OMC timeline and plan. I feel somewhat on shaky ground with the answers, but figured I'd post them so I can be corrected once and for all.

  1. Is the plan really to use the current OMC setup to make a homodyne measurement? 
    1. I'm not sure where on the timeline the new OMC and BHD switchover are relative to each other. I have been imagining doing the swap to BHD before having a new OMC.
  2. I thought the current OMC resurrection plan was to do DC readout and not homodyne?
    1. I think the OMC resurrection plan is leading to DC readout, but when we switch over to BHD we'll be able to operate at the dark fringe. Is that right?
  3. Is it really possible to use our single OMC to clean both the LO and dark port beams? Isn't this the whole raging debate for A+?
    1. My understanding is yes, with the LO and DP in orthogonal polarizations. It's not clear to me why we expect to be able to do this while there is a debate for A+, perhaps our requirements are weaker. It is something I should calculate, I'll talk to Koji.
  4. A layout diagram would be really useful.
    1. Attached now.
  5. Where in the priority list does this come in?
    1. I am a leaf in the wind. I think this comes well after we have the OMC resurrected, we just want to get a sense for what parts we need so we can order them before the fiscal year closes.
Quote:

That seems fine, I wasn't thinking of that beam. in that case could we just have a PBS directly behind MMT2 and send both beams to the same OMMT?

Alternatively we can move OM5 and the beam path OMPO-OMMTSM towards -y, then put the LO-OMMT parallel to the existing OMMT but displaced in +x... we'd have to move the existing OMC and BHD towards +x as well. 

Quote:

Can we use the leakage beam from MMT2 on the OMC table as the LO beam? I can't find the spec for this optic, but the leakage beam was clearly visible on an IR card even with the IMC locked with 100 mW input power so presumably there's enough light there, and this is a cavity transmission beam which presumably has some HOM content filtered out.

 

Attachment 1: BHD_layout.pdf
BHD_layout.pdf
  14157   Mon Aug 13 11:44:32 2018 gautamUpdateComputer Scripts / ProgramsPatch updates on nodus

Larry W said that some security issues were flagged on nodus. So I ran

sudo yum upgrade --exclude=elog-3.1.3-2.el7.x86_64

on nodus. The exclude flag is because there were some conflicts related to that particular package. Hopefully this has fixed the problem. It's been a while since the last update, which was in January of this year.

controls@nodus|~> sudo yum history
Loaded plugins: langpacks
ID     | Command line             | Date and time    | Action(s)      | Altered
-------------------------------------------------------------------------------
    29 | upgrade --exclude=elog-3 | 2018-08-13 11:36 | E, I, U        |  136 EE
    28 | install yum-utils        | 2018-08-13 11:31 | Update         |    1   
    27 | install nmap             | 2018-06-29 01:57 | Install        |    2   
    26 | install grace            | 2018-05-31 16:52 | Install        |   11   
    25 | install https://dl.fedor | 2018-05-31 16:51 | Install        |    1   
    24 | install perl-Digest-SHA1 | 2018-05-31 15:34 | Install        |    1   
    23 | install python-devel     | 2018-01-13 15:33 | Install        |    1   
    22 | install gcc              | 2018-01-13 15:32 | Install        |    6   
    21 | install git              | 2018-01-12 18:11 | Install        |    4   
    20 | update                   | 2018-01-12 18:01 | I, U           |   39   
    19 | install motif            | 2018-01-05 17:35 | Install        |    3   
    18 | install sendmail sendmai | 2017-12-03 05:11 | Install        |    6   
    17 | install vim              | 2017-11-21 18:12 | Install        |    3   
    16 | reinstall mod_dav_svn    | 2017-11-21 17:40 | Reinstall      |    1   
    15 | install mod_dav_svn      | 2017-11-21 17:39 | Install        |    1   
    14 | install subversion       | 2017-11-21 15:36 | Install        |    2   
    13 | -y install php           | 2017-11-20 22:15 | Install        |    4   
    12 | install links            | 2017-11-20 19:10 | Install        |    2   
    11 | install openssl098e.i686 | 2017-11-18 18:28 | Install        |    1   
    10 | install openssl-libs.i68 | 2017-11-18 18:26 | Install        |   11   
history list
  14156   Mon Aug 13 09:56:23 2018 SteveUpdateSUSETMX trip follow-up

Here is an other big one

Quote:

A brief follow-up on this since we discussed this at the meeting yesterday: the attached DV screenshot shows the full 2k data for a period of 2 seconds starting just before the watchdog tripped. It is clear that the timescale of the glitch in the UL channel is much faster (~50 ms) compared to the (presumably mechanical) timescale seen in the other channels of ~250 ms, with the step also being much smaller (a few counts as opposed to the few thousand counts seen in the UL channel, and I guess 1 OSEM count ~ 1 um). All this supports the hypothesis that the problem is electrical and not mechanical (i.e. I think we can rule out the Acromag sending a glitchy signal to the coil and kicking the optic). The watchdog itself gets tripped because the tripping condition is the RMS of the shadow sensor outputs, which presumably exceeds the set threshold when UL glitches by a few thousand counts.

 

Attachment 1: ETMXglitch.png
ETMXglitch.png
Attachment 2: ETMXgltch.png
ETMXgltch.png
  14155   Sun Aug 12 10:59:34 2018 aaronConfigurationUpgradeParts list for BHD

That seems fine, I wasn't thinking of that beam. in that case could we just have a PBS directly behind MMT2 and send both beams to the same OMMT?

Alternatively we can move OM5 and the beam path OMPO-OMMTSM towards -y, then put the LO-OMMT parallel to the existing OMMT but displaced in +x... we'd have to move the existing OMC and BHD towards +x as well. 

Quote:

Can we use the leakage beam from MMT2 on the OMC table as the LO beam? I can't find the spec for this optic, but the leakage beam was clearly visible on an IR card even with the IMC locked with 100 mW input power so presumably there's enough light there, and this is a cavity transmission beam which presumably has some HOM content filtered out.

  14154   Fri Aug 10 16:43:50 2018 gautamConfigurationUpgradeParts list for BHD

Can we use the leakage beam from MMT2 on the OMC table as the LO beam? I can't find the spec for this optic, but the leakage beam was clearly visible on an IR card even with the IMC locked with 100 mW input power so presumably there's enough light there, and this is a cavity transmission beam which presumably has some HOM content filtered out.

Quote:

My current thought is to use the MC reflection, the beam that heads from MC1 to MCR1, as the LO beam

  14153   Fri Aug 10 11:29:39 2018 aaronConfigurationUpgradeParts list for BHD

I've started putting together a list of things we'll need to buy to do BHD readout. I'm still messing around with more detailed optics layouts, but wanted to get a list started here so people can let me know if I'm missing any big, obvious categories of goods.

My current plan makes minimal changes to the signal path going to the OMC, and tries to just get the LO beam into the OMC with minimal optics. I'm not thinking of any of the optics as suspended, and it requires several reflections of the LO beam, so probably this is not an excellent configuration, but it's a start for getting the parts list:

  1. My current thought is to use the MC reflection, the beam that heads from MC1 to MCR1, as the LO beam
    1. From MCR1, send the LO to a BS that directs it into an MMT, oriented along x (and lets us keep the MC refl PO)
    2. After the two MMT optics, the beam will be traveling along -x, and can be directed to a mirror that sends it towards -y to two steering mirrors that send it along -x then +x near the top of the table (perpendicular to the signal MMT.
    3. Then, send it to a PBS, which replaces the mirror directly after the signal MMT. This is where it combines 
  2. Beam is then sent to the steering mirrors to bring it into the OMC
  3. In parallel, the signal beam is going through the same path it has now, including the pickoff beam, with the one change that we need a HWP somewhere before the PBS, and the PBS replaces the mirror directly after the MMT (and needs to move a bit closer to have the beam properly directed)

I started making a layout of this scheme, but it's probably not going to work so I'm going to make a quick layout of this more major modification instead:

  1. Both the MCR beam and the AS beam come in about parallel. Send each to a PO mirror.
  2. The PO mirror directs both beams into parallel MMT aligned along x
  3. From the MMT, each is directed to a pair of steering mirrors located where the OMC MMT is located now
  4. From the steering mirrors go to the PBS that combines the signal and LO
  5. Then to two more steering mirrors to get into the OMC, which may be moved towards +x
  6. From the OMC go to the BHD PBS

What we need

Optics

  • HWP for just before the LO combines with the signal
  • HWP for just before the signal combines with the LO (is this necessary?)
  • PBS to replace OM5 (combines the LO and the signal)
  • Two MMT optics
  • Two piezo-driven TT optics for steering the LO to the PBS
  • One additional non-piezo optic for between the LOMMT and the LO-TTs
  • One additional BS to get the LO into the MMT (and to let us still have the PO)
  • -1 optic—we pick up one mirror that we replace with the PBS

Optomechanics

  • 2x HWP mounts
  • 1x PBS mount
  • 2x mounts for piezo-driven TT
  • 2x MMT optic mounts—I didn’t take a close enough look at these during the vent to know what we need here
  • 2x mounts for ordinary optics
  • 9x clamps for optics mounts (maybe fewer if some are on blocks)
  • 9x posts for optics mounts

Electronics

  • Additional TT driver
  • HV supply for the new TTs
  • Are the HWP actively controlled? We might need something to drive those?
  • Do we have enough DAC/ADC channels?

Questions

These are mostly just miscellaneous

  1. What of these optics need to be suspended? If we need suspensions on all of the LO optics, including the MMT, I’m not sure we’re going to be able to fit all of this on the table as I envision it…..
  2. What if anything can we put out of vacuum (HWP for example)?
  3. Do we actually need two MMT?
  14152   Fri Aug 10 01:10:56 2018 gautamUpdateLSCSome vertex locking restored

For the first time after the whirlwind vent, I managed to lock the PRMI.

  • First, I did POX/POY locking, dither aligned the arms to maximize TRX and TRY.
  • Next, I misaligned the ETM and tested the Michelson locking
    • Since we've lost ~70% of power on the AS55 PD, I set the whitening gain for AS55 I and Q channels to +6dB (old value was 0dB).
    • worked alright. In this new config, the peak-to-peak Michelson fringe count is ~80 cts, while I reported ~60cts-pp a couple of months ago, so all seems good on that front.
    • But the config script in the IFOconfigure MEDM screen somehow doesn't set the AS55_Q ----> MICH_A element in the LSC input matrix anymore.
    • I edited the .snap file for this configuration to set the relevant matrix element EPICS channel to +1.0.
    • I also edited the overall loop gain for this configuration from +30 to +2 (for bright fringe, use -2 for dark fringe).
  • Feeling adventerous, I decided to try PRMI in the carrier resonant tuning (to be clear, PRCL on REFL11_I, MICH on AS55_Q).
    • Finding the REFL spot on the camera took a while since the PRM has been macroscopically misaligned for the mode-scanning
    • Went out to the table and centered the REFL beam onto REFL11 and REFL55 PDs - didn't need much tweaking, which is a good sign, since we shouldn't have screwed anything up on the symmetric side by any of the vent activities.
    • Restored PRMI locking using the IFOconfigure MEDM screen - lock caught almost immediately.
    • Ran the dither alignment servos for MICH and PRCL - BS needed a bit of encouragement to make the dark spot dark, but POP has been pretty stable over ~15mins.
    • I didn't take any loop transfer functions, to do.

I don't have the energy to make a DRMI attempt tonight - but the signs are encouraging. I'd like to use the IFO in the next few days to try and recover DRMI locking. The main concern is that the optical path on the AS beam has changed by ~0.3m I estimate. So the demod phase for AS55 may need to be adjusted, but the change due to optical path length only should be ~10degrees so the DRMI locking with the old settings should still work. Perhaps we also want to scan the PRC and SRC with the phase information from the Trans/Refl transfer functions as well.


Don't want to jinx it, but the c1lsc FE models have been stable. Tomorrow, I'd like to re-enable c1cal, since it has some useful channels for NBing. Could c1daf/c1oaf which have significant amounts of custom C code be the culprits?

Attachment 1: PRMIcarrier.png
PRMIcarrier.png
  14151   Thu Aug 9 22:50:13 2018 gautamUpdateCDSAlignSoft script modified

After this work of increasing the series resistance on ETMX, there have been numerous occassions where the insufficient misalignment of ETMX has caused problems in locking vertex cavities. Today, I modified the script (located at /opt/rtcds/caltech/c1/medm/MISC/ifoalign/AlignSoft.py) to avoid such problems. The way the misalign script works is to write an offset value to the "TO_COIL" filter bank (accessed via "Output Filters" button on the Suspension master MEDM screen - not the most intuitive place to put an offset but okay). So I just increased the value of this offset from 250 counts to 2500 counts (for ETMX only). I checked that the script works, now when both ETMs are misaligned, the AS55Q signal shows a clean Michelson-like sine wave as it fringes instead of having the arm cavity PDH fringes as well yes.

Note that the svn doesn't seem to work on the newly upgraded SL7 machines: svn status gives me the following output.

svn: E155036: Please see the 'svn upgrade' command
svn: E155036: Working copy '/cvs/cds/rtcds/userapps/trunk/cds/c1/medm/MISC/ifoalign' is too old (format 10, created by Subversion 1.6)

 Is it safe to run 'svn upgrade'? Or is it time to migrate to git.ligo.org/40m/scripts?

Attachment 1: MichelsonFringing.png
MichelsonFringing.png
  14150   Thu Aug 9 12:40:14 2018 gautamUpdateSUSETMX trip follow-up

A brief follow-up on this since we discussed this at the meeting yesterday: the attached DV screenshot shows the full 2k data for a period of 2 seconds starting just before the watchdog tripped. It is clear that the timescale of the glitch in the UL channel is much faster (~50 ms) compared to the (presumably mechanical) timescale seen in the other channels of ~250 ms, with the step also being much smaller (a few counts as opposed to the few thousand counts seen in the UL channel, and I guess 1 OSEM count ~ 1 um). All this supports the hypothesis that the problem is electrical and not mechanical (i.e. I think we can rule out the Acromag sending a glitchy signal to the coil and kicking the optic). The watchdog itself gets tripped because the tripping condition is the RMS of the shadow sensor outputs, which presumably exceeds the set threshold when UL glitches by a few thousand counts.

Attachment 1: ETMXglitch.png
ETMXglitch.png
  14149   Thu Aug 9 12:31:13 2018 gautamUpdateCDSCDS status update

The model seems to have run without issues overnight. Not completely related, but the MC1 shadow sensor signals also don't show any abnormal excursions to negative values in the last 48 hours. I'm thinking about re-connecting the satellite box (but preserving the breakout setup at 1X6 for a while longer) and re-locking the IMC. I'll also start c1ass on the c1lsc frontend. I would say that the other models on c1lsc (i.e. c1oaf, c1cal, c1daf) aren't really necessary for basic IFO operation.

Quote:

As part of this slow but systematic debugging, I am turning on the c1lsc model overnight to see if the model crashes return.

  14148   Thu Aug 9 02:12:13 2018 gautamUpdateCOCSouth East or West?

Summary:

For operating the SRC in the "Signal-Recycled" tuning, the SRC macroscopic length needs to be ~4.04m (compared to the current value of ~5.399m), assuming we don't do anything fancy like change the modulation frequencies and not transmit through the IMC. We're putting together a notebook with all the calculations, but today I was thinking about what the signal extraction path should be, specifically which chamber the SRM should be in. Just noting down the thoughts I had here while they're fresh in my head, all this has to be fleshed out, maybe I'm making this out to be more of a problem than it actually is.

Details:

  • For the current modulation frequencies, if we want the reosnance conditions such that the f2 sideband is resonant in the SRC (but not f1, i.e. small Schnupp asymmetry regime) while the carrier is resonant in the arms (required for good sensing of the SRC length), the macroscopic length of the SRC needs to be changed to ~4.04m.
  • Practically, this means that the folded SRC would only have one folding mirror (SR2).
  • There is a shorter SRC length of ~1.something metres which would work, but that would involve changing the relative position between ITMs and BS (currently ~2.3m) so I reject that option for now.
  • So the SR2 would be roughly where it is right now, ~20cm from the BS.
  • The question then becomes, where do we direct the reflection from the SR2? We need an optical path length of ~1.5m from SR2. So options are 
    • ITMY table (East)
    • ITMX table (South)
    • IMC table (West)
  • Moreover, after the SRM, we have to accommodate:
    • Some kind of pickoff for in-air PDs.
    • OFI.
    • OMC MMT.
    • OMC.
  • Some kind of CBA (as of now I think going to the ITMY table is the best option):
Option Advantages Disadvantages
ITMY
  • Easy to direct beam from BS/PRM chamber to the ITMY table (i.e. we don't have to worry too much about avoiding other optics in the path etc).
  • Ease of access to chamber, ease of working in there.
  • ITMY table probably has the most room to work out an OFI + OMC MMT + OMC solution.
  • AS beam extraction to air will be more complicated, possibly have to do it on ITMY optical table.
  • Not sure if the ITMY table can accommodate all of the output optics subsystems I listed above.
  • Routing the LO beam to this table would be tricky I guess.
ITMX
  • Routing the LO beam for homodyne detection is probably easiest in this chamber.
  • Allows for small AoI on folding mirror, reducing the impact of astigmatism.
  • Pain to work in this chamber because of IMC tube.
  • Steering beam from SR2 to ITMX table means threading the needle between PRM and PR3 possibly.
IMC
  • Probably allows the use of (almost) the entire existing OMC chamber for the output optics (OFI, OMC MMT, OMC).
  • IMC table is crowded (2 SOS towers, several steering optics for the input beam, input faraday).
  • Not sure what is the performance of the seismic isolation stacks on these tables vs the larger optical tables.
  • Painful to work in these smaller chambers.
  14147   Wed Aug 8 23:06:59 2018 gautamUpdateSUSAnother low noise bias path idea

Today while Rich Abbott was here, Koji and I had a brief discussion with him about the HV amplifier idea for the coil driver bias path. He gave us some useful tips, perhaps most useful being a topology that he used and tested for an aLIGO ITM ESD driver which we can adapt to our application. It uses a PA95 high voltage amplifier which differs from the PA91 mainly in the output voltage range (up to 900V for the former, "only" 400V for the former. He agrees with the overall design idea of 

  • Having a LN opamp with the HV amp inside the feedback loop for better voltage noise at low frequencies.
  • Having a passive RC network at the output of the HV amp to filter out noise at high frequencies.

He also gave some useful suggestions like 

  • Using the front panel of the box that as a heatsink for the HV amps.
  • Testing the stability of the nested opamp loop by "pinging" the output of the opamp with some pulses from a function generator and monitoring the response to this perturbation on a scope.

I am going to work on making a prototype version of this box for 5 channels that we can test with ETMX. I have been told that the coupling from side coil to longitudinal motion is of the order of 1/30, in which case maybe we only need 4 channels.

  14146   Wed Aug 8 23:03:42 2018 gautamUpdateCDSc1lsc model started

As part of this slow but systematic debugging, I am turning on the c1lsc model overnight to see if the model crashes return.

  14145   Wed Aug 8 20:56:11 2018 KojiUpdatePSLEOM measuement preparation

Rich and I worked on the EOM measurement. After the measurement, the setup was reverted to the nominal state

  • AUX PLL mixer was restored to ZAD-6
  • The PLL gain was restored to 3.10
  • The main PSL marconi is connected to the freq generator again. Using the beat note, I've confirmed that the modulations are applied on the beam.
  • The PSL HEPA was reduced from 100 to 30.
  14144   Tue Aug 7 23:06:30 2018 KojiUpdatePSLEOM measuement preparation

I was preparing for the aLIGO EOM measuement to be carried out tomorrow afternoon.

I did a few modifications to the PLL setup.

  • The freq mixier in the PLL setup was replaced with ZP3 (level 7) from ZAD-6
  • The PLL gain was reduced from 3.10 to 2.80 to prevent servo oscillation
  • The main PSL marconi is connected to the PLL mixier and providing fixed 200MHz 8dBm.
  • The main PSL modulation is off.

Tomorrow I am going to modulate the EOM with the AUX Marconi via an amplifier (probably)

Automated scripts (AGinit.py and AGmeas.py) are in /users/koji/scripts

I will revert the setup once the measurement is done tomorrow.

  14143   Tue Aug 7 22:28:23 2018 gautamUpdateCDSMore CDS woes

I am starting the c1x04 model (IOP) on c1lsc to see how it behaves overnight.

Well, there was apparently an immediate reaction - all the models on c1sus and c1ioo reported an ADC timeout and crashed. I'm going to reboot them and still have c1x04 IOP running, to see what happens.

[97544.431561] c1pem: ADC TIMEOUT 3 8703 63 8767
[97544.431574] c1mcs: ADC TIMEOUT 1 8703 63 8767
[97544.431576] c1sus: ADC TIMEOUT 1 8703 63 8767
[97544.454746] c1rfm: ADC TIMEOUT 0 9033 9 8841
Quote:

Overnight, all models on c1sus and c1ioo seem to have had no stability issues, supporting the hypothesis that timing issues stem from c1lsc. Moreover, the MC1 shadow sensor readouts showed no negative values over a ~12hour period. I think we should just observe this for another day, in any case I don't think there is any urgent IFO related activity scheduled.

  14142   Tue Aug 7 11:30:46 2018 gautamUpdateCDSMore CDS woes

Overnight, all models on c1sus and c1ioo seem to have had no stability issues, supporting the hypothesis that timing issues stem from c1lsc. Moreover, the MC1 shadow sensor readouts showed no negative values over a ~12hour period. I think we should just observe this for another day, in any case I don't think there is any urgent IFO related activity scheduled.

  14141   Mon Aug 6 20:41:10 2018 aaronUpdateDAQNew DAC for the OMC

Gautam and I tested out the DAC that he installed in the latter half of last week. We confirmed that at least one of the channels is can successfully drive a sine wave (ch10, 1-indexed). We had to measure the output directly on the SCSI connector (breakout in the FE hard drive cabinet along the Y arm), since the SCSI breakout box (D080303) seems not to be working (wiring diagram in Gautam's elog from his SURF years).

I added some DAC channels to our c1omc model:
PZT1_PIT
PZT1_YAW
PZT2_PIT
PZT2_YAQ
 
And determined that when we go to use the ADC, we will initially want the following channels (even these are probably unnecessary for the very first scans):
TRANS_PD1
TRANS_PD2
REFL_PD
DVMDC (drive voltage monitor, DC level)
DVMAC ("", AC level, only needed if we dither the length)
 
I attach a screenshot of the model, and a picture of where the whitening/dewhitening boards should go in the rack.
Attachment 1: OMCDACmdl.png
OMCDACmdl.png
  14140   Mon Aug 6 19:49:09 2018 gautamUpdateCDSMore CDS woes

I've left the c1lsc frontend shutdown for now, to see if c1sus and c1ioo can survive without any problems overnight. In parallel, we are going to try and debug the MC1 OSEM Sensor problem - the idea will be to disable the bias voltage to the OSEM LEDs, and see if the readback channels still go below zero, this would be a clear indication that the problem is in the readback transimpedance stage and not the LED. Per the schematic, this can be done by simply disconnecting the two D-sub connectors going to the vacuum flange (this is the configuration in which we usually use the sat box tester kit for example). Attachment #1 shows the current setup at the PD readout board end. The dark DC count (i.e. with the OSEM LEDs off) is ~150 cts, while the nominal level is ~1000 cts, so perhaps this is already indicative of something being broken but let's observe overnight.

Attachment 1: IMG_7106.JPG
IMG_7106.JPG
  14139   Mon Aug 6 14:38:38 2018 gautamUpdateCDSMore CDS woes

Stability was short-lived it seems. When I came in this morning, all models on c1lsc were dead already, and now c1sus is also dead (Attachment #1). Moreover, MC1 shadow sensors failed for a brief period again this afternoon (Attachment #2). I'm going to wait for some CDS experts to take a look at this since any fix I effect seems to be short-lived. For the MC1 shadow sensors, I wonder if the Trillium box (and associated Sorensen) failure somehow damaged the MC1 shadow sensor/coil driver electronics.

Quote:
 

Let's see how stable this configuration is. Onto some locking now...

Attachment 1: CDScrash.png
CDScrash.png
Attachment 2: MC1failures.png
MC1failures.png
  14138   Mon Aug 6 09:42:10 2018 KojiSummaryComputersTransition of the main NFS disk on chiara

Follow up:

- At least it was confirmed that the local backup (4TB->2TB) is regularly running every morning.

- The 2TB disk was used up to 95%. To ease the size of the remaining space, I have further compressed the burt snapshot folders. (~2016). This released another 150GB. The 2TB is currently used up to  87%.

Prev

Filesystem      1K-blocks       Used  Available Use% Mounted on
/dev/sdc1      3845709644 1731391748 1918967020  48% /home/cds
/dev/sdd1      2113786796 1886162780  120249888  95% /media/40mBackup

Now

Filesystem      1K-blocks       Used  Available Use% Mounted on
/dev/sdc1      3845709644 1731706744 1918652024  48% /home/cds
/dev/sdd1      2113786796 1728124828  278287840  87% /media/40mBackup

 

  14137   Mon Aug 6 09:34:02 2018 SteveUpdateVACRGA scan at day 20

 

 

Attachment 1: pd81d20.png
pd81d20.png
  14136   Mon Aug 6 00:26:21 2018 gautamUpdateCDSMore CDS woes

I spent most of today fighting various CDS errors.

  • I rebooted c1lsc around 3pm, my goal was to try and do some vertex locking and figure out what the implications were of having only ~30% power we used to have at the AS port.
  • Shortly afterwards (~4pm), c1lsc crashed.
  • Using the reboot script, I was able to bring everything back up. But the DC lights on c1sus models were all red, and a 0x4000 error was being reported.
  • This error is indicative of some timing issue, but all the usual tricks (reboot vertex FEs in various order, restart the mx_streams etc) didn't clear this error.
  • I checked the Tempus GPS unit, that didn't report any obvious problems (i.e. front display was showing the correct UTC time).
  • Finally, I decided to shut down all watchdogs, soft reboot all the FEs, soft reboot FB, power cycle all expansion chassis.
  • This seems to have done the trick - I'm leaving c1oaf disabled for now.
  • The remaining red indicators are due to c1dnn and c1oaf being disabled.

Let's see how stable this configuration is. Onto some locking now...

Attachment 1: CDSoverview.png
CDSoverview.png
  14135   Sun Aug 5 15:43:50 2018 gautamUpdateSUSAnother low noise bias path idea

OK, how about this:

  • Attachment #1 shows the proposed schematic.
    • It consists of a second order section with Gain x10 to map the +/-10V DC range of the DAC to +/- 100V DC such that we preserve roughly the same amount of DC actuation range.
    • Corner frequency of the SOS is set to ~0.7 Hz. In hindsight, maybe this is more aggressive than necessary, we can tune this.
    • DC gain is 20 dB (typo in the text where I say the DC gain is x15, though we could go with this option as well I think if we want a larger series resistance).
    • A first order passive low-pass stage is added to filter out the voltage noise of the PA91, which dominates the output voltage noise (next bullet).
  • Attachment #2 shows the transfer function from input to output
    • The two traces compare having just a single SOS filtering stage vs the current topology of having two SOS stages.
    • The passive output RC network is necessary in either case to filter the voltage noise of the PA91 OpAmp.
    • For the DAC noise, I just assumed a flat noise level of 5 \mu V / \sqrt{\mathrm{Hz}}, I don't actually know what this is for the Acromag DACs.
  • Attachments #3 shows a breakdown of the top 5 noise contributions.
    • The PA91 datasheet doesn't give current noise information so I just assumed 1 fA / \sqrt{\mathrm{Hz}}, which was what was used for the PA85 in the existing opamp.lib file.
    • The voltage noise is modelled as 4.5 \sqrt{1+\frac{80}{f}} nV / \sqrt{\mathrm{Hz}}, which seems to line up okay with the plot on Pg4 of the datasheet.
    • So the model suggests we will be dominated by the voltage noise of the PA91.
  • Attachment #4 translates the noise into current noise seen by the actuator.
    • I add the Johnson noise contribution of the series resistance for this path, which is assumed to be 10 k \Omega.
    • For comparison, I add the filtered DAC noise contribution, and Johnson noise of the proposed series resistance in the fast path.
    • For the bias path, we are dominated by the Johnson noise of the series resistor from ~60 Hz upwards.
    • It's not quite fair to say that the Johnson noise of the resistance in the fast path dominates, the quadrature sum of fast and bais paths will be ~1.2 times of the former alone. 
    • Bottom line: we will be in the regime of total current noise of ~2.2 pA/rtHz, where I think Kevin's modeling suggests we can see some squeezing.

The question still remains of how to combine the fast and bias paths in this proposed scheme. I think the following approach works for prototyping at least:

  • Remove the series resistance on the existing coil driver boards' bias path, hence isolating this from the coil.
  • Route the DB15 output connector from the coil driver board (which is now just the fast actuation signals) into a sub-sattelite box housing the bias path electronics.
  • Sum the two signals as it is done now, by simply having a conductor (PCB trace) merge the two paths after their respective series resistances.

In the longer term, perhaps the Satellite Box revamp can accommodate a bias voltage summation connector.

Quote:

Bah! Too complex.


I have neglected many practical concerns. Some things that come to mind:

  1. Is it necessary to protect the upstream DAC from some potential failure of the PA91 in which the high voltage appears at the input?
  2. What is the correct OpAmp for this purpose? This chart on Apex's page suggests that PA15, PA85, PA91 and PA98 are all comparable in terms of drive capability, and the spec sheets don't suggest any dramatic differences. Some LIGO circuits use PA85, some use PA90, but I can't find any that use PA91. Perhaps Rana/Koji can comment about this.
  3. What kind of protection is necessary for the PA91 power?
  4. What is the correct way to do heat management? Presumably we need heatsinks, and in fact, there is a variant of the packaging style that has "formed" legs, which from what I can figure out, allow the heat sink plane on the PA91 to be parallel to the PCB surface. But I think the heat-sink wisdom suggests vertical fins are the most efficient (not sure if this holds if the PCB is inside a box though). What about the PCB itself? Are some kind of special traces needed?
  5. Can we use the current-limiting resistor feature on the PA91? The datasheet seems to advice against it for G>10 configurations, which is what we need, although our requirement is only at DC so I don't know if that table is applicable to this circuit.
  6. Are 3W resistors sufficient? I think we require only 10mA maximum current to preserve the current actuation range, so 100 V * 10mA = 1W, so 3W leaves some safety margin.
  7. All capacitors should be rated for 500 V per the datasheet.  
Attachment 1: HV_Bias_schematic.pdf
HV_Bias_schematic.pdf
Attachment 2: TF.pdf
TF.pdf
Attachment 3: bias.pdf
bias.pdf
Attachment 4: HVbias_currentNoise.pdf
HVbias_currentNoise.pdf
  14134   Sun Aug 5 13:45:00 2018 gautamUpdateSUSETMX tripped

Independent from the problems the vertex machine has been having (I think, unless it's something happening over the shared memory network), I noticed on Friday that the ETMX watchdog was tripped. Today, once again, the ETMX watchdog was tripped. There is no evidence of any abnormal seismic activity around that time, and anyways, none of the other watchdogs tripped. Attachment #1 shows that this happened ~838am PT today morning. Attachment #2 shows the 2k sensor data around the time of the trip. If the latter is to be believed, there was a big impulse in the UL shadow sensor signal which may have triggered the trip. I'll squish cables and see if that helps - Steve and I did work at the EX electronics rack (1X9) on Friday but this problem precedes our working there...

Attachment 1: ETMX_tripped.png
ETMX_tripped.png
Attachment 2: ETMX_tripped_zoom.png
ETMX_tripped_zoom.png
  14133   Sun Aug 5 13:28:43 2018 gautamUpdateCDSc1lsc flaky

Since the lab-wide computer shutdown last Wednesday, all the realtime models running on c1lsc have been flaky. The error is always the same:

[58477.149254] c1cal: ADC TIMEOUT 0 10963 19 11027
[58477.149254] c1daf: ADC TIMEOUT 0 10963 19 11027
[58477.149254] c1ass: ADC TIMEOUT 0 10963 19 11027
[58477.149254] c1oaf: ADC TIMEOUT 0 10963 19 11027
[58477.149254] c1lsc: ADC TIMEOUT 0 10963 19 11027
[58478.148001] c1x04: timeout 0 1000000 
[58479.148017] c1x04: timeout 1 1000000 
[58479.148017] c1x04: exiting from fe_code()

This has happened at least 4 times since Wednesday. The reboot script makes recovery easier, but doing it once in 2 days is getting annoying, especially since we are running many things (e.g. ASS) in custom configurations which have to be reloaded each time. I wonder why the problem persists even though I've power-cycled the expansion chassis? I want to try and do some IFO characterization today so I'm going to run the reboot script again but I'll get in touch with J Hanks to see if he has any insight (I don't think there are any logfiles on the FEs anyways that I'll wipe out by doing a reboot). I wonder if this problem is connected to DuoTone? But if so, why is c1lsc the only FE with this problem? c1sus also does not have the DuoTone system set up correctly...

The last time this happened, the problem apparently fixed itself so I still don't have any insight as to what is causing the problem in the first place frown. Maybe I'll try disabling c1oaf since that's the configuration we've been running in for a few weeks.

  14132   Fri Aug 3 19:02:11 2018 gautamUpdateASSX arm ASS recovery

[koji, gautam]

After I effected the series resistance change for ETMX, the X arm ASS didn't work (i.e. IR transmission would degrade if the servo was run). Today, we succeeded in recovering a functional ASS servo yes.

So both arms have working dither alignment servos now. But remember that the Y arm ASS gains have been set for locking the Y arm with MC2 as the actuator, not ETMY.

Details:

  • Koji pointed out that the demodulated signals from the ETM dither are only used to center the spot on the ETM, and that we should first run the servo with existing settings with the ETM pitch and yaw spot centering loops disabled.
    • This improved TRX level from ~0.8 to 1.1
  • Next, we tried increasing the LO amplitudes by x5 to account for the reduced actuation of the dither on ETMX
    • We then re-enabled the two loops that were earlier disabled.
    • This resulted in TRX degrading very quickly.
  • So we decided to try going back to the nominal LO gains, and reducing the gain of the two ETM spot centering loops.
    • This did the trick, TRX went from 1.1 --> ~1.23, which is the nominal maximum pre-vent value.
  • The snap file used to recover the correct settings to run the dither alignment servos have been updated, the old one has been backed up with today's datestamp.

We then tried to maximize GTRX using the PZT mirrors, but were only successful in reaching a maximum of 0.41. The value I remember from before the vent was 0.5, and indeed, with the IR alignment not quite optimized before we began this work, I saw GTRX of 0.48. But the IR dither servo signals indicate that the cavity axis may have shifted (spot position on the ITM, which is uncontrolled, seems to have drifred significantly, the Pitch signal doesn't stay on the StripTool scale anymore). So we may have to double check that the transmitted beam isn't falling off the GTRX DC PD.

  14131   Fri Aug 3 18:54:58 2018 gautamUpdateSUSGlitchy MC1

The wall StripTool indicated that the IMC wasn't too happy when I came in today. Specifically:

  • MC1 watchdog was tripped.
  • Even in the tripped state, MC REFL spot on the camera showed spot motion that was too large to be explained as normal seismic driven motion (i.e. with local damping supposedly disabled).
  • Strange excursions were observed in the MC1 shadow sensor signal levels as well, see Attachment #1 - negative values don't make any sense for this readout.

The last time this happened, it was due to the Sorensens not spitting out the correct voltages. This time, there were no indications on the Sorensens that anything was funky. So I just disabled the MCautolocker and figured I'd debug later in the evening.

However, around 5pm, the shadow sensor values looked nominal again, and when I re-enabled the local damping, the MC REFL spot suggested that the local damping was working just fine. I re-enabled the MCautolocker, MC re-locked almost immediately. To re-iterate, I did nothing to the electronics inside the VEA. Anyways, this enabled us to work on the X arm ASS (next elog).

Attachment 1: MC1_sensorAnomaly.png
MC1_sensorAnomaly.png
  14130   Fri Aug 3 16:27:40 2018 ranaUpdateSUSLow noise bias path idea

Bah! Too complex.

  14129   Fri Aug 3 15:53:25 2018 gautamUpdateSUSLow noise bias path idea

Summary:

The idea we are going with to push the coil driver noise contribution down is to simply increase the series resistance between the coil driver board output and the OSEM coil. But there are two paths, one for fast actuation and one that provides a DC current for global alignment. I think the simplest way to reduce the noise contribution of the latter, while preserving reasonable actuation range, is to implement a precision DC high-voltage source. A candidate that I pulled off an LT application note is shown in Attachment #1.

Requirements:

  • The series resistance in the bias path should be 10 k\Omega, such that the noise from this stage is dominated by the Johnson noise of said resistor, and hence, the current noise contribution is negligible compared to the series resistance in the fast actuation path (4.5 k\Omega).
  • Since we only really need this for the test masses, what actuation range do we want?
    • Currently, ETMY has a series resistance of 400\Omega and has a pitch DC bias voltage of -4 V. 
    • This corresponds to 10 mA of DC current.
    • To drive this current through 10 k\Omega, we need 100 V. 
    • I'm assuming we can manually correct for yaw misalignments such that 10mA of DC current will be sufficient for any sort of corrective alignment.
    • So +/- 120 V DC should be sufficient.
  • The current noise of this stage should be negligible at 100 Hz. 
    • The noise of the transistors and the HV supply should be suppressed by the feedback loop and so shouldn't be a significant contribution (I'll model to confirm).
    • The input noise of the LT1055 is ~20nV/rtHz at 100 Hz, while the Johnson noise of 10 k\Omega is ~13nV/rtHz so maybe the low-passing needs to be tuned, but I think if it comes to it, we can implement a passive RC network at the output to achieve additional filtering.
  • To implement this circuit, we need +/- 125V DC. 
    • At EX and EY, we have a KEPCO HV supply meant to be used for the Green Steering PZTs. 
    • I'm not sure if these can do bipolar outputs, if not, for temporary testing, we can transport the unit at EY to EX.

If all this seems reasonable, I'd like to prototype this circuit and test it with ETMX, which already has the high series resistance for the fast path. So I will ask Steve to order the OpAmp and transistors.

Attachment 1: LT1055_precOpAmp.pdf
LT1055_precOpAmp.pdf
  14128   Fri Aug 3 14:35:56 2018 gautamSummaryElectronicsEX AUX electronics power restored

Steve and I restored the power to the EX AUX electronics rack. The power strip on the lowest shelf of the AUX rack now goes to another power strip laid out vertically along the NW corner of 1X9. The EX green locks to the arm just fine now.

  14127   Thu Aug 2 23:09:25 2018 ranaSummaryComputersX Green "Mystery" solved

I'm going to guess that this was me: I was disconnecting some octopus power strip nonsense down there (in particular illuminators and cameras), so I might have turned off the AUX rack by mistake.

Quote:

I walked down to the X end and found that the entire AUX laser electronics rack isn't getting any power. There was no elog about this.

I couldn't find any free points in the power strip where I think all this stuff was plugged in so I'm going to hold off on resurrecting this until tomorrow when I'll work with Steve.

Quote:

The X arm green does not stay locked to the cavity - the alignment looks fine, and the green flashes are strong, but the lock does not hold. This shouldn't be directly connected to anything we did today since the Green PDH servo is entirely analog.

  14126   Thu Aug 2 20:54:18 2018 gautamSummaryComputersc1omc model looks stable

Actually, c1lsc had crashed again sometime last night so I had to reboot everything this morning. I used the reboot script again, but I increased the sleep time between trying to start up the models again so that I could walk into the VEA and power cycle the c1lsc expansion chassis, as this kind of frequent model crash has been fixed by doing so in the past. Sure enough, there have been no issues since I rebooted everything at ~1030 in the morning. 

The c1omc model itself has been stable as well, though of course, there is nothing in there at the moment. I may do a check of the newly installed DAC tomorrow just to see that we can put out a sine wave.

Steve has ordered the D-sub cabling that will allow us to route signals between AA/AI boards in 1X1/1X2 to the HV PZT electronics in the OMC rack. Things look setup for a measurement next week. Aaron will post a block diagram + photoz of what box goes where in the electronics racks.

  14125   Thu Aug 2 20:47:29 2018 gautamSummaryElectronicsX Green "Mystery" solved

I walked down to the X end and found that the entire AUX laser electronics rack isn't getting any power. There was no elog about this.

I couldn't find any free points in the power strip where I think all this stuff was plugged in so I'm going to hold off on resurrecting this until tomorrow when I'll work with Steve.

Quote:

The X arm green does not stay locked to the cavity - the alignment looks fine, and the green flashes are strong, but the lock does not hold. This shouldn't be directly connected to anything we did today since the Green PDH servo is entirely analog.

  14124   Thu Aug 2 16:30:08 2018 SteveUpdateTreasuretime capsule location

I 've just found this time capsule note from Nov. 26, 2000 by Kip Thorne:  LIGO will discover gravitational waves by Dec.31, 2007

Quote:

   Beautifully Done

   Chirp

  what is next?

Atm 3, Ron Drever could not celebrate with us because of health issues.

 

 

Attachment 1: time_capsule.JPG
time_capsule.JPG
ELOG V3.1.3-